# Differentially Private Covariate Balancing Causal Inference

### v1

@@ Abstract
Differential privacy is the leading mathematical framework for privacy protection, providing a probabilistic guarantee that safeguards individuals’ private information when publishing statistics from a dataset. This guarantee is achieved by applying a randomized algorithm to the original data, which introduces unique challenges in data analysis by distorting inherent patterns. In particular, causal inference using observational data in privacy-sensitive contexts is challenging because it requires covariate balance between treatment groups, yet checking the true covariates is prohibited to prevent leakage of sensitive information. In this article, we present a differentially private two-stage covariate balancing weighting estimator to infer causal effects from observational data. Our algorithm produces both point and interval estimators with statistical guarantees, such as consistency and rate optimality, under a given privacy budget.
@@ 1 Introduction
In light of the unique challenges associated with robust causal analysis in privacy-sensitive contexts, this article introduces a privacy-preserving, covariate-balancing causal inference methodology. Our work contributes to the current body of literature in several ways. First, we propose a two-stage, differentially private algorithm for estimating a wide range of causal effects from observational data. By employing covariate balancing weighting schemes, our method effectively addresses misspecification in the propensity score model, while preserving privacy. Second, we establish the differential privacy guarantees of our algorithm within a specified privacy budget. Additionally, we derive the asymptotic properties of our estimator, demonstrating consistency, rate optimality, and asymptotic covariate balance. We also provide asymptotically valid confidence intervals for the estimated causal effects. Third, through comprehensive simulation studies across various privacy budgets and sample sizes, we evaluate the performance of our methodology and compare it against existing DP causal inference approaches. The results indicate that our methodology exhibits robust performance under both correctly specified and misspecified propensity score models. Finally, we apply our methodology to real-world data from a job training program evaluation, successfully recovering the non-private estimates and achieving satisfactory covariate balance. This application underscores the practical utility and effectiveness of our approach in privacy-sensitive contexts. All technical proofs of theorems and lemmas are provided in the supplementary material.
@@ 2.1 Notation and Causal Estimands
[…]
@@
where $h(x)$ is a known function of the covariates. By choosing different forms of the function $h(\cdot)$, the WATE can capture various estimands of causal effects: $h(x)=1$ for ATE, $h(x)=e(x)$ for the ATE on the treated (ATT), $h(x)=1-e(x)$ for the ATE on the control (ATC).
@@ 2.2 Covariate Balancing Propensity Score Estimation
with $\hat{\theta}_{\alpha,\beta}$ being estimated with corresponding score functions for the specific values of $\alpha$ and $\beta$. Table 1 summarizes the discussion in this section about the correspondence of estimands, sample weighting functions, and score functions. Our methodology starts with choosing the causal quantity of interest and corresponding values for $\alpha$ and $\beta$, through which the score function is automatically defined.
@@
$\alpha$ $\beta$ $\tau_{\alpha,\beta}$ $w(x,1)$ $w(x,0)$ $S(e,1)$ $S(e,0)$1 -1 ATE $\frac{1}{e(x)}$ $\frac{1}{1-e(x)}$ $\log\frac{e}{1-e}-\frac{1}{e}$ $\log\frac{1-e}{e}-\frac{1}{1-e}$1 0 ATC $\frac{1-e(x)}{e(x)}$ $1$ $\log\frac{1-e}{e}$ $-\frac{1}{e}$ 0 -1 ATT $\frac{1}{e(x)}$ $\frac{1}{1-e(x)}$ $\log\frac{e}{1-e}$ $-\frac{1}{1-e}$ 0 0 ATO $1-e(x)$ $e(x)$ $\log e$ $\log(1-e)$
@@ 2.3 Differential Privacy
Another important concept in DP is sensitivity. The probabilistic guarantee of DP mechanisms is often achieved by adding random noise to the statistics of interest. Importantly, the noise must be scaled proportionally to the sensitivity of the statistics, which measures the worst-case magnitude by which the statistics may change between two adjacent databases. Formally, the $\ell_{1}$-sensitivity of a function $f$: $\mathcal{D}\to\mathbb{R}^{k}$ is $\Delta_{f}=\sup_{D,D^{\prime}\in\mathcal{D}_{n}}||f(D)-f(D^{\prime})||_{1}$. One of the most commonly used DP mechanisms is the Laplace mechanism, which adds noise to a function of interest.
@@
Among the methods to achieve DP, the exponential mechanism (McSherry and Talwar, 2007) is another popular mechanism for its flexibility and adaptability across various statistical analyses. This mechanism is especially useful for problems involving the minimization of an objective function $\ell_{n}(\theta,D)$ for $D\in\mathcal{D}_{n}$, which encompasses a wide range of statistical tasks. However, as Awan et al. (2019) demonstrated, the noise added by the exponential mechanism can be significant, sometimes exceeding that of other mechanisms, thereby reducing its utility. Another popular mechanism that involves minimizing an objective function is the output/objective perturbation (Chaudhuri et al., 2011), which Lee et al. (2019) adopted for the propensity score estimation. A limitation of these methods for our purposes is their requirement of regularization, which introduces bias that can remain even asymptotically.
Reimherr and Awan (2019) introduced the K-Norm Gradient Mechanism (KNG) as an alternative to objective perturbation by altering the typical use of the exponential mechanism.
@@
[…]
@@
One condition for the KNG to achieve $O_{p}(n^{-1})$ error is that the loss function is strongly convex. Without the strong convexity (only with convexity), the KNG is still a valid $\epsilon$-DP mechanism, which is an additional advantage over other mechanisms that require strong convexity (e.g., objective perturbation), but may have error greater than $O_{p}(n^{-1})$. The negative score function $-S\{e_{\theta}(x),z\}$, which we use as the loss function in this study, is a convex function for $z\in\{0,1\}$ and $-1\leq\alpha,\beta\leq 1$, and $S$ is strongly convex for all $-1\leq\alpha,\beta\leq 1$ except for $\alpha=-1,\beta=0$ and $\alpha=0,\beta=-1$ (Zhao, 2019, Theorem 3.2).
@@ 3.2 Minimax Risk Lower Bound for WATE Estimation
First, we discuss a lower bound for the WATE estimation problem under the central DP model, where a data curator has access to individual data and then applies a differential privacy mechanism and releases the privatized outputs (e.g., summary statistics) to an untrusted data analyst. According to Barber and Duchi (2014), the minimax lower bound of the mean squared error for 1-dimensional mean estimation under the central model is $O[\{n^{-1}+(\epsilon n)^{-2}\}]$. We let $\mathcal{M}_{\epsilon}$ denote the set of all privacy mechanisms that satisfy $\epsilon$-DP. Suppose $\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{n}$ are drawn according to some distribution $P\in\mathcal{P}$, where $\mathcal{P}$ denotes a class of distributions on the sample space of potential outcomes. Also, We define an estimator $\hat{\tau}$ for WATE $\tau$ as a measurable function that maps inputs to a real value, that is, $\hat{\tau}:\Omega^{n}\to\mathbb{R}$, where $\Omega$ generally denotes the space of inputs.
@@
Lemma 1 is a simple modification of Barber and Duchi (2014, Proposition 2), implying that if a DP WATE estimation procedure achieves the minimax lower bound $O\{n^{-1}+(\epsilon n)^{-2}\}$, then it is minimax optimal among all $\epsilon$-DP procedures. Failing to match the bound does not necessarily imply the method is suboptimal. It still means the lower bound may not be tight.
@@ 3.3 Private Propensity Score Estimation
$\alpha$ $\beta$ $\tau_{\alpha,\beta}$ $\Delta_{\theta,\alpha,\beta}$ $\Delta_{V,\alpha,\beta}$1 -1 ATE $2C_{\phi}/\eta$ $(2n\eta)^{-1}$1 0 ATC $2C_{\phi}(1-\eta)/\eta$ $(2n\eta^{2})^{-1}$ 0 -1 ATT $2C_{\phi}(1-\eta)/\eta$ $(2n\eta^{2})^{-1}$ 0 0 ATO $2C_{\phi}(1-\eta)$ $(2n\eta^{2}(1-\eta))^{-1}$
We first consider estimating the propensity score in a private manner via the KNG mechanism with $K=2$. We consider privatizing the parameter $\hat{\theta}_{n}$ obtained by solving (2). Given a privacy budget $\epsilon$, we draw $\tilde{\theta}^{(1)}_{\alpha,\beta}$ from the density:
@@
Since we adopt Assumption 1, we truncate $e_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i})$ to $[\eta,1-\eta]$, which is needed to learn the sensitivity in the next section. $\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}$ is obtained by plugging in $e_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i})$ for the weighting function in Table 1. Note that the estimator (11) does not satisfy $\epsilon$-DP and is vulnerable to adversarial attacks because it contains private information. Thus, we need an additional privatization step, which we develop in the following section. However, before moving to the second stage, we present the following theorem.
@@ 3.4 Private WATE Estimation
where $\nu_{j}\sim\mathrm{Lap}\{0,1/(1-p)q_{j}\epsilon\eta\}$ and $10<q_{j}<$ 1 for $j=1,\ldots,4$, and $\sum_{j=1}^{4}q_{j}=1$ and the $\ell_{1}$-sensitivity of each component is $1/\eta$. The quantities $q_{j}$ represent the customizable privacy budget for each component of (11). Algorithm 1 summarizes the process for obtaining the privatized point estimator $\tilde{\tau}^{(2)}_{\alpha,\beta}$. Proposition 3 establishes the privacy guarantee of Algorithm 1.
Input: Database $D=\{(X_{i},Z_{i},Y_{i})\}_{i=1}^{n}$, causal estimands of interest in Table 1, the corresponding score function $S_{n,\alpha,\beta}$ and sensitivity $\Delta_{\theta,\alpha,\beta}$, positivity bound $\eta$, privacy budget $\epsilon$, fraction of privacy budget allocation parameters $p$ and $q_{j}$ for $j=1,\ldots,4$.
Output: Differentially private estimator $\tilde{\tau}^{(2)}_{\alpha,\beta}$
@@
    First Stage: Privatization of Propensity Score
    1. (a)
      Using the privacy-aware rejection sampler (Awan and Rao, 2024), draw the private propensity score parameter $\tilde{\theta}^{(1)}_{\alpha,\beta}$ from: $f(\theta)\propto\exp\left\{-\frac{p\epsilon}{2\Delta_{\theta,\alpha,\beta}}\|\nabla S_{n,\alpha,\beta}(\theta;D)\|_{2}\right\}$.
    2. (b)
@@
    3. (c)
      Truncate $e_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i})$ to $[\eta,1-\eta]$.
@@
    Second Stage: Release of the Final Privatized Estimator
    1. (a)
    2. (b)
      Output the following estimator: $\tilde{\tau}^{(2)}_{\alpha,\beta}=\frac{\sum_{i=1}^{n}Z_{i}\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},1)Y_{i}+\nu_{1}}{\sum_{i=1}^{n}Z_{i}\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},1)+\nu_{2}}-\frac{\sum_{i=1}^{n}(1-Z_{i})\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},0)Y_{i}+\nu_{3}}{\sum_{i=1}^{n}(1-Z_{i})\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},0)+\nu_{4}}$, where $\nu_{j}\sim\text{Lap}\left\{0,\frac{1}{(1-p)q_{j}\epsilon\eta}\right\}$ for $j=1,\dots,4$.
Algorithm 1 Differentially Private Covariate Balancing Estimator.

### v2

@@ Abstract
Differential privacy is the leading mathematical framework for privacy protection, providing a probabilistic guarantee that safeguards individuals’ private information when publishing statistics from a dataset. This guarantee is achieved by applying a randomized algorithm to the original data, which introduces unique challenges in data analysis by distorting inherent patterns. In particular, causal inference using observational data while preserving privacy is challenging because it requires a good covariate balance between treatment groups, but checking the covariate balance is challenging when privacy is a primary concern because releasing any statistic (e.g., t-statistic) for the balance check compromises privacy. Additionally, the performance of the privatized estimator critically depends on the choice of differential privacy mechanisms, and it remains unexplored how privacy-protecting causal inference could be conducted with observational data. In this article, we present a differentially private two-stage covariate balancing weighting estimator to infer causal effects from observational data. Our algorithm produces both point and interval estimators with statistical guarantees, such as consistency and rate optimality, under a given privacy budget.
@@ 1 Introduction
In light of the unique challenges associated with robust causal analysis in privacy-sensitive contexts, this article introduces a privacy-preserving, covariate-balancing causal inference methodology. Our work contributes to the current body of literature in several ways. First, we discuss privatization strategies and the selection of privacy mechanisms to achieve desirable asymptotic properties, such as consistency and rate optimality. We then propose a two-stage privatization algorithm within a unified framework for estimating a wide range of causal effects from observational data. Leveraging the covariate-balancing scoring rules (CBSR) of Zhao (2019), we obtain weights that are robust to propensity-score misspecification in privacy-sensitive settings. This choice of balancing framework is pivotal for selecting a compatible privacy mechanism, as its convex, differentiable objective pairs naturally with DP methods that privatize gradients. We instantiate privacy with the K-Norm Gradient (KNG) mechanism (Reimherr and Awan, 2019), which releases a noise-perturbed score and solves the resulting estimating equations, thereby preserving CBSR’s balance properties while ensuring privacy. By contrast, generic mechanisms such as the exponential mechanism (McSherry and Talwar, 2007) and objective/output perturbation (Chaudhuri et al., 2011) do not, in general, maintain the required balance constraints or utility guarantees for weighting, and thus may fail to deliver the desired properties. Additionally, with appropriate choices of the covariate balancing framework and privacy mechanism, our privatized estimator exhibits favorable asymptotic properties, such as consistency, rate optimality, and asymptotic covariate balance, while preserving privacy. We also provide asymptotically valid confidence intervals for the estimated causal effects. Through comprehensive simulation studies across various privacy budgets and sample sizes, we evaluate the performance of our methodology and compare it against existing DP causal inference approaches. The results indicate that our methodology exhibits robust performance under both correctly specified and misspecified propensity score models. Finally, we apply our methodology to real-world data from a job training program evaluation, successfully recovering the non-private estimates and achieving satisfactory covariate balance. This application underscores the practical utility and effectiveness of our approach in privacy-sensitive contexts. All technical proofs of theorems and lemmas are provided in the supplementary material.
The rest of the paper is organized as follows. Section 2 presents the preliminaries for the differential privacy and covariate balancing causal inference framework. Section 3 provides our alrorithm and proved its privacy guarantees and statistical properties. Section 4 provides simulation studies for validating our methodology developed in the previous sections, and Section 5 provides an application of our methodologies to real-world data of a job training program. Section 6 concludes with some final discussion.
@@ 2.1 Notation and Causal Estimands
We define the average treatment effect (ATE) as $\tau=\mathrm{E}\{Y(1)-Y(0)\}$. Assuming the existence of the marginal density function for the covariates $X$, denoted as $f(x)$, with respect to a base measure $\mu$, we can write the ATE as $\tau=\int\tau(x)f(x)dx$, where $\tau(x)=\mathrm{E}\{Y(1)-Y(0)\mid X=x\}$. The ATE is widely valued due to its ability to provide interpretable estimates of treatment effects on the whole population. However, in practice, the exclusive focus on the ATE is a significant limitation for several reasons. First, the ATE may reflect the effect of an intervention that is practically infeasible or impossible to apply to every unit in the study population. Second, the available data may not accurately capture the characteristics of the intended population of interest. In such cases, conventional analysis methods may fail to yield an estimate of the average treatment effect for the appropriate target population. Finally, researchers frequently exclude extreme or atypical units from their analysis, resulting in estimates that pertain only to a subpopulation within the original target population. A recent body of literature suggests focusing on subpopulations exhibiting sufficient covariate overlap between treatment groups. Li et al. (2018) introduced a class of balancing weights designed to balance the distributions of covariates between comparison groups for any predetermined target population. They consider the target population density, expressed as the product of the marginal density $f(x)$ and a predefined function $h(x)$ of the covariate $x$. Within this framework, a general class of estimands is defined as the weighted average treatment effect (WATE) over the target population:
@@
where $h(x)$ is a known function of the covariates. By choosing different forms of the function $h(\cdot)$, the WATE can capture various estimands of causal effects: $h(x)=1$ for ATE, $h(x)=e(x)$ for the ATE on the treated (ATT), h(x)=1e(x)h(x)=1-e(x) for the ATE on the control (ATC).
@@ 2.2 Covariate Balancing Propensity Score Estimation
with $\hat{\theta}_{\alpha,\beta}$ being estimated with corresponding score functions for the specific values of $\alpha$ and $\beta$. Table 1 summarizes the discussion in this section about the correspondence of estimands, sample weighting functions, and score functions. Note that our methodology starts with choosing the causal quantity of interest and corresponding values for $\alpha$ and $\beta$, through which the score function is automatically defined.
@@
$\alpha$ $\beta$ $\tau_{\alpha,\beta}$ $w(x,1)$ $w(x,0)$ $S(e,1)$ $S(e,0)$ -1 -1 ATE $\frac{1}{e(x)}$ $\frac{1}{1-e(x)}$ $\log\frac{e}{1-e}-\frac{1}{e}$ $\log\frac{1-e}{e}-\frac{1}{1-e}$ -1 0 ATC $\frac{1-e(x)}{e(x)}$ 11 $\log\frac{1-e}{e}$ $\frac{1}{e}$ 0 -1 ATT $\frac{1}{e(x)}$ $\frac{1}{1-e(x)}$ $\log\frac{e}{1-e}$ $\frac{1}{1-e}$ 0 0 ATO 1e(x)1-e(x) $e(x)$ $\log e$ $\log(1-e)$
Remark (Synthesis with privacy mechanism).
Among covariate-balancing propensity score methods, entropy balancing (Hainmueller, 2012) is widely used. It reweights units by solving a maximum-entropy program that enforces prespecified covariate moments to match across groups, yielding a balanced pseudo-population under ignorability. Chan et al. (2015) propose empirical balancing calibration weighting, which constructs calibration weights by minimizing a divergence from uniform subject to linear balance constraints on chosen basis functions. However, both approaches are ill-suited for privacy-preserving weighting: standard privacy mechanisms require bounded losses or gradients to calibrate noise, whereas the entropy and calibration objectives generally violate this boundedness. By contrast, the CSBR framework satisfies the required boundedness and is therefore compatible with private weighting.
@@ 2.3 Differential Privacy
[…]
@@
Reimherr and Awan (2019) introduced the K-Norm Gradient Mechanism (KNG). This mechanism is especially useful for problems involving the minimization of an objective function through which the parameters of interest are obtained.
@@
Let $\Theta\subset\mathbb{R}^{d}$ be a convex set, $\|\cdot\|_{K}$ be a norm on $\mathbb{R}^{d}$, and $\nu$ be the Lebesgue measure on $\Theta$. Let $\{\ell_{n}(\theta;D):\Theta\rightarrow\mathbb{R}\mid D\in\mathcal{D}_{n}\}$ be a collection of measurable functions whose gradient is defined almost everywhere. We say that this collection has sensitivity $\Delta:\Theta\rightarrow\mathbb{R}_{+}$, if $\|\nabla\ell_{n}(\theta;D)-\nabla\ell_{n}(\theta;D^{\prime})\|_{K}\leq\Delta(\theta)<\infty$, for all adjacent $D,D^{\prime}\in\mathcal{D}_{n}$ and $\theta$. If $\int_{\Theta}\exp\left\{-\frac{1}{\Delta(\theta)}\|\nabla\ell_{n}(\theta;D)\|_{K}\right\}d\nu(\theta)<\infty$ for all $D\in\mathcal{D}_{n}$, then the collection of probability measures $\{\mu_{D}\mid D\in\mathcal{D}_{n}\}$ with densities (with respect to $\nu$) given by
@@
One condition for the KNG to achieve $O_{p}(n^{-1})$ error is that the loss function is strongly convex. Without the strong convexity (only with convexity), the KNG is still a valid $\epsilon$-DP mechanism, which is an additional advantage over other mechanisms that require strong convexity (e.g., objective perturbation), but may have error greater than $O_{p}(n^{-1})$. The negative score function $S\{e_{\theta}(x),z\}$, which we use as the loss function in this study, is a convex function for $z\in\{0,1\}$ and $1\leq\alpha,\beta\leq 1$, and $S$ is strongly convex for all $1\leq\alpha,\beta\leq 1$ except for $\alpha=-1,\beta=0$ and $\alpha=0,\beta=-1$ (Zhao, 2019, Theorem 3.2).
Remark (Choice of privacy mechanism).
Among the methods to achieve DP, the exponential mechanism (McSherry and Talwar, 2007) is a popular mechanism for its flexibility and adaptability across various statistical analyses. This mechanism is especially useful for problems involving the minimization of an objective function $\ell_{n}(\theta,D)$ for $D\in\mathcal{D}_{n}$, which encompasses a wide range of statistical tasks. The exponential mechanism releases an estimate $\tilde{\theta}$ based on the density: $f(\theta)\propto\exp\{-c_{0}\ell_{n}(\theta,D)\},$ where $c_{0}$ is a constant determined by the sensitivity of $\ell_{n}$ and the desired level of privacy. However, as Awan et al. (2019) demonstrated, the noise added by the exponential mechanism can be significant, sometimes exceeding that of other mechanisms, thereby reducing its utility. Another popular mechanism that involves minimizing an objective function is the output/objective perturbation (Chaudhuri et al., 2011), which Lee et al. (2019) adopted for the propensity score estimation. A limitation of these methods for our purposes is their requirement of regularization, which introduces bias that can remain even asymptotically and breaks the covariate balance and leads to inconsistent estimates. Additionally, the regularization parameter is typically determined via cross-validation. Each iteration of cross-validation increases the risk of information leak, thus, it is desirable to define the objective function for the propensity score without regularization to maintain the balance.
@@ 3.2 Minimax Risk Lower Bound for WATE Estimation
First, we discuss a lower bound for the WATE estimation problem under the central DP model, where a data curator has access to individual data and then applies a differential privacy mechanism and releases the privatized outputs (e.g., summary statistics) to an untrusted data analyst. According to Barber and Duchi (2014), the minimax lower bound of the mean squared error for 1-dimensional mean estimation under the central model is $O[\{n^{-1}+(\epsilon n)^{-2}\}]$. We let $\mathcal{M}_{\epsilon}$ denote the set of all privacy mechanisms that satisfy $\epsilon$-DP. Suppose $\{(X_{i},Y_{i},Z_{i})\}_{i=1}^{n}$ are drawn according to some distribution $P\in\mathcal{P}$, where $\mathcal{P}$ denotes a class of distributions on the sample space of covariates, potential outcomes and treatment assignment variables. Also, we define an estimator $\hat{\tau}$ for WATE $\tau$ as a measurable function that maps inputs to a real value, that is, $\hat{\tau}:\Omega^{n}\to\mathbb{R}$, where $\Omega$ generally denotes the space of inputs.
@@
Lemma 1 is a simple modification of Barber and Duchi (2014, Proposition 2), implying that if a DP WATE estimation procedure achieves the minimax lower bound $O\{n^{-1}+(\epsilon n)^{-2}\}$, then it is minimax optimal among all $\epsilon$-DP procedures. Failing to match the bound does not necessarily imply the method is suboptimal as it may be the case that the lower bound is simply not tight.
@@ 3.3 Private Propensity Score Estimation
$\alpha$ $\beta$ $\tau_{\alpha,\beta}$ $\Delta_{\theta,\alpha,\beta}$ $\Delta_{V,\alpha,\beta}$ -1 -1 ATE $2C_{\phi}/\eta$ $(2n\eta)^{-1}$ -1 0 ATC $2C_{\phi}(1-\eta)/\eta$ $(2n\eta^{2})^{-1}$ 0 -1 ATT $2C_{\phi}(1-\eta)/\eta$ $(2n\eta^{2})^{-1}$ 0 0 ATO $2C_{\phi}(1-\eta)$ $(2n\eta^{2}(1-\eta))^{-1}$
We first consider estimating the propensity score in a private manner via the KNG mechanism with $\lVert\cdot\rVert_{K}=\lvert\cdot\rVert_{2}$. We consider privatizing the parameter $\hat{\theta}_{n}$ obtained by solving (2). Given a privacy budget $\epsilon$, we draw $\tilde{\theta}^{(1)}_{\alpha,\beta}$ from the density:
@@
Since we adopt Assumption 1, we truncate $e_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i})$ to $\eta,1-\eta]$, which is needed to learn the sensitivity in the next section. $\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}$ is obtained by plugging in $e_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i})$ for the weighting function in Table 1. Note that the estimator (11) does not satisfy $\epsilon$-DP and is vulnerable to adversarial attacks because it contains private information. Thus, we need an additional privatization step, which we develop in the following section. Prior to that, we first present the following theorem.
@@ 3.4 Private WATE Estimation
where $\nu_{j}\sim\mathrm{Lap}\{0,1/(1-p)q_{j}\epsilon\eta\}$ and $0<q_{j}<1$ for $j=1,\ldots,4$, and $\sum_{j=1}^{4}q_{j}=1$ and the $\ell_{1}$-sensitivity of each component is $1/\eta$. The quantities $q_{j}$ represent the customizable privacy budget for each component of (11). Algorithm 1 summarizes the process for obtaining the privatized point estimator $\tilde{\tau}^{(2)}_{\alpha,\beta}$. Proposition 3 establishes the privacy guarantee of Algorithm 1.
Algorithm 1 Differentially Private Covariate Balancing Estimator
0: Database $D=\{(X_{i},Z_{i},Y_{i})\}_{i=1}^{n}$, causal estimands of interest in Table 1, the corresponding score function $S_{n,\alpha,\beta}$ and sensitivity $\Delta_{\theta,\alpha,\beta}$, positivity bound $\eta$, privacy budget $\epsilon$, fraction of privacy budget allocation parameters $p$ and $q_{j}$ for $j=1,\ldots,4$.
0: Differentially private estimator $\tilde{\tau}^{(2)}_{\alpha,\beta}$.
1:First Stage: Privatization of Propensity Score
@@
    Using the privacy-aware rejection sampler (Awan and Rao, 2024), draw the private propensity score parameter $\tilde{\theta}^{(1)}_{\alpha,\beta}$ from:
    $f(\theta)\propto\exp\left\{-\frac{p\epsilon}{2\Delta_{\theta,\alpha,\beta}}\|\nabla S_{n,\alpha,\beta}(\theta;D)\|_{2}\right\}.$
  2. 2.
@@
  3. 3.
    Truncate $e_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i})$ to $\eta,1-\eta]$.
2:Second Stage: Release of the Final Privatized Estimator
  1. 1.
@@
    Output the following estimator:
    $\tilde{\tau}^{(2)}_{\alpha,\beta}=\frac{\sum_{i=1}^{n}Z_{i}\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},1)Y_{i}+\nu_{1}}{\sum_{i=1}^{n}Z_{i}\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},1)+\nu_{2}}-\frac{\sum_{i=1}^{n}(1-Z_{i})\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},0)Y_{i}+\nu_{3}}{\sum_{i=1}^{n}(1-Z_{i})\tilde{w}_{\tilde{\theta}^{(1)}_{\alpha,\beta}}(X_{i},0)+\nu_{4}},$
    where $\nu_{j}\sim\text{Lap}\left\{0,\frac{1}{(1-p)q_{j}\epsilon\eta}\right\}$ for $j=1,\dots,4$.